If the price is set at $72 per unit, what production quota is needed to make sure there are no shortages or surpluses?


Considering the price support and quota, calculate:

the consumer surplus

the producer surplus

deadweight loss


1
Expert's answer
2021-10-18T11:29:43-0400

Solution:

A production quota of 30 units equal to the new quantity demanded is needed to make sure there are no shortages or surpluses.


Consumer surplus = "\frac{1}{2} \times30 \times (90-72 = 270"


Producer surplus = "\frac{1}{2} \times30 \times (72-15) = 855"


Deadweight loss = "\frac{1}{2} \times(50-30) \times(72-45) = 270"